# STAR: A Simple Training-free Approach for Recommendations using Large Language Models

### v1

@@ 3.3.1. Rank schema
We present three main strategies for ranking: (1) Point-wise evaluates each item $x\in I_{k}$ independently, based on the user sequence $S_{u}$, to determine how likely it is that user $u$ will interact with item $x$. If two items receive the same score, their rank follows the initial order from $I_{k}$; (2) Pair-wise evaluates the preference between two items $x_{i},x_{j}\in I_{k}$ based on the user sequence $S_{u}$. We adopt a sliding window approach, starting from the items with the lowest retrieval score at the bottom of the list (qin-etal-2024-large, ). The LLM compares and swaps adjacent pairs, while iteratively stepping the comparison window one element at a time. (3) List-wise evaluates the preference among multiple items $x_{i},\dots,x_{i+w}\in I_{k}$ based on the user sequence $S_{u}$. This method also uses a sliding window approach, with a window size $w$ and a stride $d$ to move the window across the list, refining the ranking as it passes (sun2023chatgpt, ). In this setup, pair-wise is a special case of list-wise with $w=2$ and $d=1$.
3.3.2. Item information
@@
We represent the metadata (e.g., Item ID, title, category, etc.) for each item in the user sequence $s_{j}\in S_{u}$ and each candidate item to be ranked $x\in I_{k}$ as JSON format in the input prompt. Additionally, we incorporate two more types of information that can help the reasoning capabilities of the LLM: (1) Popularity is calculated as the number of users who have interacted with the item $x$, simply by counting the occurrences in the training data. This popularity value is then included in the prompt for both the items in the user sequence $s_{j}\in S_{u}$ and the candidate item to be ranked $x\in I_{k}$ as “Number of users who bought this item: ###”; (2) Co-occurrence is calculated as the number of users who have interacted with both item $x$ and item $s_{j}\in S_{u}$. The resulting value is then included for candidate items $x\in I_{k}$ as “Number of users who bought both this item and item $s_{j}$: ###”.
4. Experimental Setup
@@
We evaluate the performance using public 2014 Amazon review datasets (mcauley2015image, ; he2016ups, ). Specifically, we select the Beauty, Toys and Games, and Sports and Outdoors categories, as these have been used in previous studies (geng2022recommendation, ; tan2024idgenrec, ) and provide data points for comparison (see Table 1). We follow the same data processing steps as in prior work, filtering out users and items with fewer than five interactions, maintaining consistent baseline settings.
@@
Table 1. Dataset statistics. Density is the percentage of actual user-item interactions amongst all possible user-item interactions.
@@
We compare our model with following supervised trained models: (1) KNN is a user-based collaborative filtering method that finds the top 10 most similar users to a given user and averages their ratings to score a specific item; (2) Caser uses convolution neural networks to model user interests (tang2018personalized, ); (3) HGN uses hierarchical gating networks to capture both long and short-term user behaviors (ma2019hierarchical, ); (4) GRU4Rec employs GRU to model user action sequences (hidasi2015session, ); (5) FDSA uses a self-attentive model to learn feature transition patterns (zhang2019feature, ); (6) SASRec uses a self-attention mechanism to capture item correlations within a user’s action sequence (kang2018self, ); (7) BERT4Rec applies a masked language modeling (MLM) objective for bi-directional sequential recommendation (sun2019bert4rec, ); (8) $\bf{S^{3}}$-Rec extends beyond the MLM objective by pre-training with four self-supervised objectives to learn better item representations (zhou2020s3, ). (9) P5 fine-tunes a pre-trained LM for use in multi-task recommendation systems by generating tokens based on randomly assigned item IDs  (geng2022recommendation, ); (10) TIGER also fine-tunes LMs to predict item IDs directly, but these IDs are semantic, meaning they are learned based on the content of the items (rajput2024recommender, ); and (11) IDGenRec goes further by extending semantic IDs to textual IDs, enriching the IDs with more detailed information (tan2024idgenrec, ).
@@
Unless otherwise specified, we use Gecko text-embedding-004 (lee2024gecko, ) [^1] https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text-embeddings-api to collect LLM embeddings for retrieval, and we use gemini-1.5-flash [^2] https://deepmind.google/technologies/gemini/flash/ for LLM-based ranking. All API calls were completed as of September 1, 2024.
@@
Caser (tang2018personalized, ) 0.021 0.013 0.035 0.018 0.017 0.011 0.027 0.014 0.012 0.007 0.019 0.010
HGN (ma2019hierarchical, ) 0.033 0.021 0.051 0.027 0.032 0.022 0.050 0.028 0.019 0.012 0.031 0.016
GRU4Rec (hidasi2015session, ) 0.016 0.010 0.028 0.014 0.010 0.006 0.018 0.008 0.013 0.009 0.020 0.011
BERT4Rec (sun2019bert4rec, ) 0.020 0.012 0.035 0.017 0.012 0.007 0.020 0.010 0.012 0.008 0.019 0.010
FDSA (zhang2019feature, ) 0.027 0.016 0.041 0.021 0.023 0.014 0.038 0.019 0.018 0.012 0.029 0.016
SASRec (kang2018self, ) 0.039 0.025 0.061 0.032 0.046 0.031 0.068 0.037 0.023 0.015 0.035 0.019
$\bf{S^{3}}$-Rec (zhou2020s3, ) 0.039 0.024 0.065 0.033 0.044 0.029 0.070 0.038 0.025 0.016 0.039 0.020
P5 (zhou2020s3, ) 0.016 0.011 0.025 0.014 0.007 0.005 0.012 0.007 0.006 0.004 0.010 0.005
TIGER (geng2022recommendation, ) 0.045 0.032 0.065 0.038 0.052 0.037 0.071 0.043 0.026 0.018 0.040 0.023
IDGenRec (tan2024idgenrec, ) 0.062 0.049 0.081 0.054 0.066 0.048 0.087 0.055 0.043 0.033 0.057 0.037
@@
Table 2. Performance (Hits@K, NDCG@K) comparison among supervised models, and STAR retrieval & ranking pipeline. The first group in the table represents supervised models; The second group shows the retrieval pipeline with parameters set to an exponential decay rate of $\lambda=0.7$, history length of $l=3$, and a weight factor of $a=0.5$; The third group consists of ranking pipeline which use gemini-1.5-flash. The best model for each dataset is shown in bold, and the second best is underlined.
5. Experimental Results
@@
In this section, we present a multifaceted performance analysis of our retrieval (Section 5.1) and ranking pipeline (Section 5.2).
5.1. Retrieval Pipeline
@@
The second group of Table 2 presents the performance of our retrieval framework. STAR-Retrieval alone achieves the best or second-best results compared to all baselines and fine-tuned methods. There is a significant improvement across all metrics for Toys and Games, ranging from +26.50% to +35.3%. In Beauty, all metrics besides NDCG@5 (-1.2%) are improved from a range of +6.1% to +20.0%. In Sports and Outdoors, the results are second best to IDGenRec, trailing from a range of -19.6% to -5.57%. Furthermore, we conduct more studies to answer the following research questions:
Q1. How do semantic and collaborative information affect predictions?
To evaluate the impact of semantic information ($R_{\text{S}}$) and collaborative information ($R_{\text{C}}$) on retrieval performance, we conducted an analysis by adjusting the weighting factor $a$ between these two components. As shown in the left panel of Figure 4, the optimal performance is achieved from a range of 0.5 to 0.6. We choose 0.5 for other experiments as a simple, equal weighting. Furthermore, we see that when we only use one of the two components ($a=0.0$ or $a=1.0$) then the results are significantly worse than the combination. This validates our hypothesis that it is better to include both semantic and collaborative information and that having both types of information is additive to the overall system.
Q2. How does the number of user history items $l$ and recency factor $\lambda$ affect predictions?
To assess the impact of the number of user history items $l$ and the recency factor $\lambda$ on retrieval performance, we conducted an analysis by varying both parameters. As depicted in the right panel of Figure 4, we see that having more user history items can help up to a certain point ($l=3$), but after that performance can be impacted. The middle panel of Figure 4 shows that setting an appropriate recency factor ($\lambda=0.7$) to prioritize recent history items is more effective than not applying any recency factor ($\lambda=1$). However, using a very small recenty factor can discount past items too much, which is similar to only using 1 user history item.
Refer to captionRefer to captionRefer to caption
@@
Figure 4. Retrieval performance (Hits@50) with different weighting factor $a$ between $R_{\text{S}}$ and $R_{\text{C}}$ (left), recency factor $\lambda$ (middle), number of history $l$ (right). The shaded regions show the best range of parameters. We choose $a=0.5$, $\lambda=0.7$, and $l=3$ in our experiments.
Q3. Does incorporating more user feedback improve results?
Rating Beauty Toys and Games Sports and Outdoors
@@
Table 3. Retrieval performance (Hits@10, NDCG@10) comparison between w/ and w/o rating.
In Equation  1, we propose using the user ratings to help score items. Here we test how helpful it is to provide this user feedback information. We compare the results when we score items using rating $r$ versus when we exclude this rating (by setting $r=1$ for all items). Surprisingly, ignoring the rating information consistently produced better results (See Table 3). This is likely due to a mismatch between tasks, where our analysis focuses on predicting next item interaction and we are not concerned about how a user will rate the next item. In our main evaluation and the rest of the analysis, we ignore user ratings (by setting all ratings to 1). This finding is also quite practical, because in real recommendation systems most user-item interactions do not have ratings.
Q4. Are some LLM embedding models better than others?
Figure 5. Retrieval performance (Hits@50) comparison by embedding APIs for Toys and Games. The models text-embedding-004, text-embedding-3-small, and text-embedding-3-large each have a maximum dimension of 768, 1536, and 3072, respectively.
We test different LLM embedding APIs with different embedding size to understand the impact of LLM embeddings capturing item semantic similarity . Figure 5 illustrates the performance differences between the Gecko text-embedding-004 (lee2024gecko, ), OpenAI text-embedding-3-small, and text-embedding-3-large models [^3] https://platform.openai.com/docs/guides/embeddings/. Overall, higher-dimensional embeddings and larger models performed better, indicating that enhanced semantic representation capabilities lead to improved semantic relationship capture.
Q5. Are there other ways to score items?
Scoring method Beauty Toys and Games Sports and Outdoors
@@
STAR - Retrieval ($S=1.0$, $C=0.0$) 0.072 0.042 0.095 0.055 0.039 0.022
STAR - Retrieval ($S=0.5$, $C=0.5$) 0.098 0.057 0.118 0.071 0.054 0.031
Table 4. Retrieval performance (Hits@10, NDCG@10) comparison between STAR retrieval pipeline and average embedding pooling. Here, S=Semantic Information and C=Collaborative Information weightings.
Other studies use LLM embeddings of items within a user sequence ($S_{u}$) to retrieve candidate items (harte2023leveraging, ; liang2024taxonomy, ). These methods typically involve generating a “user embedding” by performing average pooling over the LLM embeddings of the items in the user sequence. The most similar item from the candidate pool is then identified by calculating the similarity between this user embedding and the LLM embedding of each candidate item. To demonstrate the effectiveness of our scoring rule, we compare its performance against the pooling approach. As shown in Table 4, our scoring rule, even without using collaborative information for a fair comparison (i.e., $S=1.0$, $C=0.0$), outperforms the pooling approach solely using embeddings. Furthermore, incorporating collaborative information further enhances the performance (i.e., $S=0.5$, $C=0.5$).
5.2. Ranking Pipeline
The third group of Table 2 highlights the performance of the ranking framework, which improves upon the retrieval stage results. Pair-wise ranking improves all metrics over STAR-Retrieval performance by +1.7% to +7.9%. This further improves the results over other baselines for Beauty and Toys and Games, while closing the gap on IDGenRec in Sports and Outdoors. Point-wise and list-wise methods struggle to achieve similar improvements. We answer the following research questions:
Q1. Is ranking an effective approach? How do window size and stride impact performance?
@@ 5.2. Ranking Pipeline
Table 5. Ranking performance (Hits@K, NDCG@K) by window size and stride. Here we use 20 candidates from the retrieval stage. The best prompt for each dataset is shown in bold, and the second best is underlined.
Previous approaches use a selection prompt, instructing the LLM to choose the top k items in ranked order from a set of candidates (wang2023zero, ; hou2024large, ; wang-etal-2024-recmind, ). In contrast, our method uses a ranking prompt, which explicitly instructs the LLM to rank all items within the available context window. We assess the effectiveness of the ranking approach in comparison to the selection approach and a point-wise prompt. Table 5 illustrates the performance differences when varying window size and stride. The results show that ranking with a small window size (such as pair-wise or list-wise with a window size of 4) consistently outperforms the selection and pair-wise methods. Furthermore, the pair-wise and selection prompts often fail to improve performance compared to the retrieval stage, ultimately harming overall effectiveness. These findings align with previous research in document retrieval, where list-wise ranking with large window sizes or reasoning prompts requires specific task knowledge, while pair-wise ranking can be effectively performed by much smaller language models (qin-etal-2024-large, ).
Q2. How does number of candidates $k$ and number of history $l$ affect predictions?
@@
Figure 6. Pair-wise ranking performance (Hits@10) trend by different number of history $l$ and number of candidates $k$
As shown in Figure 6, varying the number of historical items included in the model, denoted as $l$, does not significantly impact prediction performance. On the other hand, increasing the number of candidate items, $k$, can be somewhat beneficial because it raises the likelihood of including the correct item among the candidates. However, the improvement in performance is minimal. Considering the high computational cost of processing a larger set of candidate items, this marginal gain may not be worthwhile.
Q3. How does additional item information in the prompt affect predictions?
Item prompt Beauty Toys and Games Sports and Outdoors
@@
Table 6. Pair-wise ranking performance (Hits@10 & NDCG@10) by varying information in the item prompt.
We investigate the impact of adding extra item information—specifically, popularity and co-occurrence data—alongside item metadata in the prompt. To assess this, we conduct an analysis by varying the types of information provided. As shown in Table 6, incorporating co-occurrence data alongside item metadata led to performance improvements in NDCG@10 ranging from +0.2% to +3.4% compared to using metadata alone. In contrast, adding popularity information did not enhance performance and, in some cases, even led to a decline. These findings suggest that introducing popularity bias does not help LLMs make better ranking decisions. This aligns with previous research indicating that simple popularity bias is ineffective in addressing recommendation problems (abdollahpouri2020multi, ; abdollahpouri2019unfairness, ; zhu2021popularity, ).
Q4. How much do LLM capabilities affect ranking performance?
Model Beauty Toys and Games Sports and Outdoors
H@10 N@10 H@10 N@10 H@10 N@10
gemini-1.5-flash 0.1008 0.0600 0.1196 0.0734 0.0564 0.0335
gemini-1.5-pro 0.1000 0.0600 0.1202 0.0754 0.0563 0.0337
gpt-4o-mini 0.1002 0.0580 0.1197 0.0722 0.0560 0.0326
gpt-4o 0.1002 0.0602 0.1205 0.0743 0.0560 0.0332
Table 7. Pair-wise ranking performance (Hits@10 & NDCG@10) comparison by different models.
To evaluate how model capabilities impact ranking performance, we compare four models: gemini-1.5-flash, gemini-1.5-pro, gpt-4o-mini, and gpt-4o. In Table 7, we observe the larger models (gemini-1.5-pro and gpt-4o) tend to outperform their smaller counterparts, although the performance differences are minimal. This finding aligns with results from other studies (qin-etal-2024-large, ), which suggest that increased model capability has limited influence on pairwise ranking tasks. Despite being more computationally expensive, pairwise ranking methods tend to be more robust than alternative approaches.
Q5. Does the order of candidates affect ranking performance?
Shuffle LLM Beauty Toys and Games Sports and Outdoors
@@
Table 8. Performance (Hits@10 & NDCG@10) comparison with random shuffling of the retrieval candidate items.
6. Conclusion

### v2

@@ 3.3.1 Rank schema
We present three main strategies for ranking: (1) Point-wise evaluates each item $x\in I_{k}$ independently, based on the user sequence $S_{u}$, to determine how likely it is that user $u$ will interact with item $x$. If two items receive the same score, their rank follows the initial order from $I_{k}$; (2) Pair-wise evaluates the preference between two items $x_{i},x_{j}\in I_{k}$ based on the user sequence $S_{u}$. We adopt a sliding window approach, starting from the items with the lowest retrieval score at the bottom of the list (Qin et al., 2024). The LLM compares and swaps adjacent pairs, while iteratively stepping the comparison window one element at a time. (3) List-wise evaluates the preference among multiple items $x_{i},\dots,x_{i+w}\in I_{k}$ based on the user sequence $S_{u}$. This method also uses a sliding window approach, with a window size $w$ and a stride $d$ to move the window across the list, refining the ranking as it passes (Sun et al., 2023). In this setup, pair-wise is a special case of list-wise with $w=2$ and $d=1$.
Figure 3: Prompt overview for the ranking pipeline. The prompt includes history items, candidate items, and instructions for the ranking strategy. Each item is represented by metadata, along with additional details such as popularity and co-occurrence, formatted in JSON. Full prompt is available in Appendix A.4.
3.3.2 Item information
@@
We represent the metadata (e.g., Item ID, title, category, etc.) for each item in the user sequence $s_{j}\in S_{u}$ and each candidate item to be ranked $x\in I_{k}$ as JSON format in the input prompt. Additionally, we incorporate two more types of information that can help the reasoning capabilities of the LLM: (1) Popularity is calculated as the number of users who have interacted with the item $x$, simply by counting the occurrences in the training data. This popularity value is then included in the prompt for both the items in the user sequence $s_{j}\in S_{u}$ and the candidate item to be ranked $x\in I_{k}$ as “Number of users who interacted with this item: ###”; (2) Co-occurrence is calculated as the number of users who have interacted with both item $x$ and item $s_{j}\in S_{u}$. The resulting value is then included for candidate items $x\in I_{k}$ as “Number of users who interacted with both this item and item $s_{j}$: ###”.
4 Experimental Setup
@@
We evaluate the performance using public 2014 Amazon review datasets (McAuley et al., 2015; He and McAuley, 2016). Specifically, we select the Beauty, Toys and Games, and Sports and Outdoors categories, as these have been used in previous studies (Geng et al., 2022; Tan et al., 2024) and provide data points for comparison (see Table 1). We follow the same data processing steps as in prior work, filtering out users and items with fewer than five interactions, maintaining consistent baseline settings.
@@
Table 1: Dataset statistics. Density is the percentage of actual user-item interactions out of all interactions.
@@
We compare our model with following supervised trained models: (1) KNN is a user-based collaborative filtering method that finds the top 10 most similar users to a given user and averages their ratings to score a specific item; (2) Caser uses convolution neural networks to model user interests (Tang and Wang, 2018); (3) HGN uses hierarchical gating networks to capture both long and short-term user behaviors (Ma et al., 2019); (4) GRU4Rec employs GRU to model user action sequences (Hidasi et al., 2015); (5) FDSA uses a self-attentive model to learn feature transition patterns (Zhang et al., 2019); (6) SASRec uses a self-attention mechanism to capture item correlations within a user’s action sequence (Kang and McAuley, 2018); (7) BERT4Rec applies a masked language modeling (MLM) objective for bi-directional sequential recommendation (Sun et al., 2019); (8) $\bf{S^{3}}$-Rec extends beyond the MLM objective by pre-training with four self-supervised objectives to learn better item representations (Zhou et al., 2020). (9) P5 fine-tunes a pre-trained LM for use in multi-task recommendation systems by generating tokens based on randomly assigned item IDs  (Geng et al., 2022); (10) TIGER also fine-tunes LMs to predict item IDs directly, but these IDs are semantic, meaning they are learned based on the content of the items (Rajput et al., 2024); and (11) IDGenRec goes further by extending semantic IDs to textual IDs, enriching the IDs with more detailed information (Tan et al., 2024).
@@
Unless otherwise specified, we use Gecko text-embedding-004 (Lee et al., 2024) [^1] https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text-embeddings-api to collect LLM embeddings for retrieval, and we use gemini-1.5-flash [^2] https://deepmind.google/technologies/gemini/flash/ for LLM-based ranking. All API calls were completed as of September 1, 2024.
@@
Caser (Tang and Wang, 2018) 0.021 0.013 0.035 0.018 0.017 0.011 0.027 0.014 0.012 0.007 0.019 0.010
HGN (Ma et al., 2019) 0.033 0.021 0.051 0.027 0.032 0.022 0.050 0.028 0.019 0.012 0.031 0.016
GRU4Rec (Hidasi et al., 2015) 0.016 0.010 0.028 0.014 0.010 0.006 0.018 0.008 0.013 0.009 0.020 0.011
BERT4Rec (Sun et al., 2019) 0.020 0.012 0.035 0.017 0.012 0.007 0.020 0.010 0.012 0.008 0.019 0.010
FDSA (Zhang et al., 2019) 0.027 0.016 0.041 0.021 0.023 0.014 0.038 0.019 0.018 0.012 0.029 0.016
SASRec (Kang and McAuley, 2018) 0.039 0.025 0.061 0.032 0.046 0.031 0.068 0.037 0.023 0.015 0.035 0.019
$\bf{S^{3}}$-Rec (Zhou et al., 2020) 0.039 0.024 0.065 0.033 0.044 0.029 0.070 0.038 0.025 0.016 0.039 0.020
P5 (Zhou et al., 2020) 0.016 0.011 0.025 0.014 0.007 0.005 0.012 0.007 0.006 0.004 0.010 0.005
TIGER (Geng et al., 2022) 0.045 0.032 0.065 0.038 0.052 0.037 0.071 0.043 0.026 0.018 0.040 0.023
IDGenRec (Tan et al., 2024) 0.062 0.049 0.081 0.054 0.066 0.048 0.087 0.055 0.043 0.033 0.057 0.037
@@
Table 2: Performance (Hits@K, NDCG@K) comparison among supervised models, and STAR retrieval & ranking pipeline. The first group in the table represents supervised models; The second group shows the retrieval pipeline with parameters set to an exponential decay rate of $\lambda=0.7$, history length of $l=3$, and a weight factor of $a=0.5$; The third group consists of ranking pipeline which use gemini-1.5-flash. The best model for each dataset is shown in bold, and the second best is underlined.
5 Experimental Results
@@
In this section, we present a performance analysis of our retrieval (§5.1) and ranking pipeline (§5.2).
5.1 Retrieval Pipeline
@@
The second group of Table 2 presents the performance of our retrieval framework. STAR-Retrieval alone achieves the best or second-best results compared to all baselines and fine-tuned methods. There is a significant improvement across all metrics for Toys and Games, ranging from +26.50% to +35.3%. In Beauty, all metrics besides NDCG@5 (-1.2%) are improved from a range of +6.1% to +20.0%. In Sports and Outdoors, the results are second best to IDGenRec, trailing from a range of -19.6% to -5.57%. Furthermore, we explore the following questions in greater detail, with additional details provided in Appendix A.1:
Impact of semantic and collaborative information.
We assess the impact of semantic ($R_{\text{S}}$) and collaborative ($R_{\text{C}}$) information by varying their weighting factor $a$. As shown in the top panel of Figure 4, the optimal performance occurs between $a=0.5$ and $0.6$, with $a=0.5$ chosen for simplicity. The far left and right sides of that figure show that results significantly degrade when using only one component ($a=0.0$ or $a=1.0$), confirming that combining both enhances retrieval effectiveness.
Impact of user history length $l$ and recency factor $\lambda$.
We analyze the effect of user history length ($l$) and recency factor ($\lambda$) on retrieval performance. As shown in the right panel of Figure 4, performance improves with more history items up to $l=3$ but declines beyond that. The middle panel indicates that a recency factor of $\lambda=0.7$ effectively prioritizes recent interactions, outperforming no recency adjustment ($\lambda=1$). However, an excessively small $\lambda$ overly discounts past items, resembling the effect of using only one history item.
Refer to captionRefer to caption
@@
Figure 4: Retrieval performance (Hits@50) with different weighting factor $a$ between $R_{\text{S}}$ and $R_{\text{C}}$ (top), recency factor $\lambda$ (bottom-left), and number of history $l$ (bottom-right). The shaded regions show the best range. $a=0.5$, $\lambda=0.7$, and $l=3$ show the best.
Impact of user rating.
Rating Beauty Toys & Games Sports & Outdoors
@@
Table 3: Retrieval (Hits@10, NDCG@10) comparing w/ and w/o incorporating user rating.
In Equation  1, we propose using the user ratings to help score items. To assess their impact, we compare results using actual ratings ($r$) versus a uniform rating ($r=1$ for all items). Surprisingly, ignoring the rating information consistently produced better results (See Table 3). This likely due to a task mismatch—our focus is on predicting interactions, not user ratings. Consequently, in our main evaluation and the rest of the analysis, we disregard ratings (by setting $r=1$ for all items), which is also practical since most real-world interactions lack explicit ratings.
Scoring rule analysis.
Scoring method Beauty Toys & Games Sports & Outdoors
@@
STAR ($S$=$1.0$, $C$=$0.0$) 0.072 0.042 0.095 0.055 0.039 0.022
STAR ($S$=$0.5$, $C$=$0.5$) 0.098 0.057 0.118 0.071 0.054 0.031
Table 4: Retrieval performance (Hits@10, NDCG@10) comparison between STAR-retrieval pipeline and average embedding pooling. $S$ representing Semantic Information and $C$ representing Collaborative Information weightings.
Prior studies (Harte et al., 2023; Liang et al., 2024) use LLM embeddings of items in a user sequence ($S_{u}$) to retrieve candidates. These methods generate a “user embedding” by averaging item embeddings, then identify the most similar candidate based on embedding similarity. To evaluate our scoring rule, we compare it against this pooling approach. As shown in Table 4, our method outperforms pooling even without collaborative information ($S=1.0$, $C=0.0$). Incorporating collaborative information ($S=0.5$, $C=0.5$) further improves performance, highlighting the advantage of our approach.
@@
Table 5: Ranking performance (Hits@K, NDCG@K) by window size and stride. Here we use 20 candidates from the retrieval stage. The best prompt for each dataset is shown in bold, and the second best is underlined.
5.2 Ranking Pipeline
The third group of Table 2 highlights the performance of the ranking framework, which improves upon the retrieval stage results. Pair-wise ranking improves all metrics over STAR-Retrieval performance by +1.7% to +7.9%. This further improves the results over other baselines for Beauty and Toys and Games, while closing the gap on IDGenRec in Sports and Outdoors. Point-wise and list-wise methods struggle to achieve similar improvements. We examine the following questions in more detail, with additional information available in Appendix A.2.
Effectiveness of ranking and impact of window size and stride.
Previous approaches use a selection prompt, instructing the LLM to choose the top-$k$ items from a candidate set (Wang and Lim, 2023; Hou et al., 2024; Wang et al., 2024b). In contrast, our method uses a ranking prompt, which explicitly instructs the LLM to rank all items within the available context window. We assess the effectiveness of the ranking approach in comparison to the selection approach and a point-wise prompt. As shown in Table 5, results show that ranking with a small window size (e.g., pair-wise or list-wise with a window of 4) consistently outperforms selection and pair-wise approaches. These findings align with prior work in document retrieval, where list-wise ranking with large windows or reasoning prompts is challenging for LLMs due to the need for task-specific knowledge. In contrast, pair-wise ranking is relatively easier and can yield better performance, even for smaller LMs (Qin et al., 2024).
Impact of candidate count ($k$) and history length ($l$).
@@ 5.2 Ranking Pipeline
Figure 5: Pair-wise ranking performance (Hits@10) trend by different number of history $l$ and number of candidates $k$
As shown in Figure 5, varying the number of historical items ($l$) has little effect on performance. In contrast, increasing the number of candidate items ($k$) slightly improves accuracy by increasing the chances of including the correct item. However, the performance gains are minimal, and given the high computational cost of processing more candidates, the trade-off may not be worthwhile.
Impact of additional item information in prompts.
Item prompt Beauty Toys & Games Sports & Outdoors
@@
Table 6: Pair-wise ranking (Hits@10 & NDCG@10) by varying information in the item prompt.
We analyze the effect of adding extra item information—specifically, popularity and co-occurrence data—alongside item metadata in prompts. As shown in Table 6, incorporating co-occurrence data improves NDCG@10 by +0.2% to +3.4% compared to using metadata alone. In contrast, adding popularity information does not enhance performance and sometimes even reduces it. This suggests that popularity bias does not help LLMs make better ranking decisions, consistent with prior research showing its ineffectiveness in recommendation tasks (Abdollahpouri and Mansoury, 2020; Abdollahpouri et al., 2019; Zhu et al., 2021).
Impact of candidate order.
Shuffle LLM Beauty Toys & Games Sports & Outdoors
@@
Table 7: Ranking (Hits@10 & NDCG@10) comparison with random shuffling of the retrieved items.
6 Conclusion